OpenStudy (help_needed_cudi):

23. Find the value of x, show all work. 5x – 9 = 16 (2 points)

OpenStudy (help_needed_cudi):

24. Find the value of t: 5t – 15 ≤ 20. Explain how you would graph the solution using a number line. (3 points) i more so need help with this question but if someone could answer both that would be awesome! *i give medals*

OpenStudy (nurali):

5x – 9 = 16 5x=16+9 5x=25 x=5 5t – 15 ≤ 20 5t≤20+15 5t≤35 t≤7 Graph: http://www.wolframalpha.com/input/?i=5t+%E2%80%93+15+%E2%89%A4+20.&dataset=

OpenStudy (anonymous):

5x – 9 = 16 16-9=7 5x/7

OpenStudy (help_needed_cudi):

thank u both!

OpenStudy (nurali):

My Pleasure.